OpenStudy (anonymous):

Please help! There is a daughter who is half the age of her mother, her mother is five years younger than her husband (father), and their ages altogether equals one-hundred (100). I have been trying to do this for almost an hour! With no such luck. Thank you!

OpenStudy (watchmath):

\(x+2x+(2x+5)=100\) What is \(x\) ?

OpenStudy (anonymous):

Oh, x is unknown.

OpenStudy (anonymous):

it would be 3x + (2x+5)=100 5x=100-5 x=19

OpenStudy (anonymous):

57 +43= 100 (Sorry, I got lost? It doesn't make sense to me.)

OpenStudy (anonymous):

let the age of daughter=x age of mother=2x age of father =2x+5 x+2x+2x+5=100 ?

OpenStudy (anonymous):